The FBI wants to determine the effectiveness of their 1010 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught.

Step 1 of 2 :

Suppose a sample of 349 suspected criminals is drawn. Of these people, 255 were not captured. Using the data, estimate the proportion of people who were caught after being on the 10 Most Wanted list. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Solution :

Given that,

n = 349

x = 255

Point estimate = sample proportion = = x / n = 255/349=0.731

=  0.731